3^{rd} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS

Krynica 29 IV – 05 V 2001

Organized by

– Institute of Mathematics of the Technical University of Lodz

Address: Al. Politechniki 11, 90-924 Łódź

Jan Kubarski (chairman), e-mail: kubarski@ck-sg.p.lodz.pl

– Institute of Mathematics of the Jagiellonian University, Cracow Address: ul. Wł. Reymonta 4, 30-059 Kraków

Robert Wolak, e-mail: wolak@im.uj.edu.pl

– Faculty of Applied Mathematics of the University Mining ang Matalurgy , Cracow

Adress: al. Mickiewicza 30, 30-059 Kraków,

Tomasz Rybicki, e-mail: tomasz@uci.agh.edu.pl

Scientific Committee: - Alekseevsky, Dmitri (Hull, GB), - Belko, Ivan (Minsk, Belarus), - Brown, Ronald (Bangor, UK), - Brzychczy, Stanisław (Cracow, Poland), - Cariñena, Jose (Zaragoza, Spain), - Grabowski, Janusz (Warsaw, Poland), - Korbaš, Julius (Bratislava, Slovakia), - Kowalski, Oldřich (Prague, Czech Republic), Kubarski, Jan (Lodz, Poland), - Mishchenko, Alexander (Moscow, Russia), - Nagy, Peter (Debrecen, Hungary), - Pradines, Jean (Toulouse, France), - Szenthe, Janos (Budapest, Hungary).

Further information available or on our WWW pages: http://im0.p.lodz.pl/konferencje/krynica2001

SECOND ANNOUNCEMENT

We have the pleasure to inform you that the planned 3rd Conference GEOMETRY AND TOPOLOGY OF MANIFOLDS has met with some interest. The list of potential participants you can find on our WWW pages.

Participants of the conference will be lodged in the Hotel ”Baśka” of the University of Mining and Metallurgy. We are going to arrange our lectures in two blocks. The first will include longer plenary lectures, which will take place in the morning. The second one, with shorter lectures will be held in the afternoon. The full day excursion in the mountains is planed.

As during the previous meetings, we would like to distribute abstracts of lectures and communications. We kindly ask to send your abstracts (LATEX in the article style or in postscript is prefered) The deadline is the 10 April 2001. We count on longer ones, even of several pages, summaries or even whole papers. They will be put on our WWW page. We plan to publish the Proceedings in the Journal ”Universitatis Iagellonicae Acta Mathematica”. If you have not sent yet the full of Registration Form including the title of your lecture please do it quickly. You can use the Registration Form on our WWW pages. Full cost of the conference (accommodation and full board) is 220 US $. The participants are requested to pay at the beginning of the conference. 29 IV is planed as the day of arriving. The bus of the Jagiellonian University Kraków-Krynica will leave Kraków about the 3 p.m. The lectures will be held 30 IV - 5 V 2001. The departure from Krynica by the bus Krynica-Kraków is 5 V in the afternoon. There is a possibility to reserve a hotel in Cracow on 28/29 IV or 5/6 V. Please inform us about it as soon as possible. The third announcement including details of the travel to Krynica will be sent in April.

Best regards, organizers:

Jan Kubarski, Robert Wolak, Tomasz Rybicki REGISTRATION FORM:

3^{rd} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS Krynica 29 IV – 05 V 2001

First and family name: Institution: Adress of the Institution: e-mail: Adress to correspondence (if different):

Date of arriving: Date of departure:

LECTURE (YES NO) Topic:

COMMUNICATION (YES NO) Topic:

The additional informations: WEGETARIAN (YES

NO)

SMOKER

(YES NO) I need a formal invitation (YES NO)

– Institute of Mathematics Technical University of Łódź Al.Politechniki 11, 90-924 Łódź,

– Institute of Mathematics of the Jagiellonian University ul. Wł. Reymonta 4, 30-059 Kraków

– Faculty of Applied Mathematics

of the University Mining ang Matalurgy, Kraków e-mail: kubarski@ck-sg.p.lodz.pl